# Rubber Wear: History, Mechanisms, and Perspectives

### v1

@@ 1 Introduction
Wear has been of concern to users for thousands of years. Even back in the Stone Age, hunters still need to consider the importance of maintaining the sharpness of their stone axes or knives. Nowadays, wear is not only a major reason for the failure of mechanical devices, but air pollution from wear particles, e.g., from tires or brakes, is one of the main threats to human health. Most particles greater than $\sim 10\ {\rm\mu m}$ (aerodynamic diameter) are deposited in the nose or throat and cannot penetrate the lower tissues of the respiratory tract. Smaller particles can enter the lungs and translocate into vital organs due to their size, causing significant human health consequences. The degree of toxicity becomes greater for smaller particles.
@@
Here we will briefly review the present understanding of different wear processes. There are several good reviews on wear. The article by P. Blau [1] focuses mainly on metals, while Muhr and Robbins [2] and Veith [3] discuss rubber wear. The history of wear, and more generally of tribology, is covered in the beautiful book by Dowson [4].
In this article, we focus on sliding wear but also briefly discuss fretting (a form of sliding wear) and erosion, where the surface of a solid is exposed to a beam of impacting hard particles. We first discuss wear mainly for metals, and later for rubber-like materials. Our main focus is on rubber wear, and we describe both the most important works of pioneers like Schallamach, Grosch, Southern and Thomas, Arnold, and Hutchings, as well as some very recent studies that attempt to include the multiscale nature of roughness on real surfaces.
Figure 1: The wear in a break. The solid bar (pin) is squeezed against the rotating cylinder (disc). If no wear of the rotating disc occur then the wear of the pin is equivalent to the disc moving towards the pin and the thickness of the removed solid material can be obtained as the difference in the circumferences between two displaced discs. The pressure is assumed to act orthogonal to the rotating disc, and the thickness of the removed material in this direction is $d(\theta)=d_{0}{\rm cos}\theta$. Hence if the wear rate is proportional to the pressure $p\sim d(\theta)\sim{\rm cos}\theta$.
2 General (mainly for metals)
One of the first scientific studies of wear was the work by K.T. Reye. He proposed 1860 that the wear is proportional to the frictional work. This is in general not true since adding a monolayer of a boundary lubricant will usually change the frictional work and the wear differently. In addition, the friction is usually due mainly to energy transfer into thermal movements and not due to wear. Still, the wear rate may be proportional to the (local) nominal contact pressure $p({\bf x})$, at least if frictional heating is unimportant, and since the frictional power often is proportional to the nominal contact pressure, $\mu pv$, if the friction coefficient us assumed constant then the local wear rate may be proportional to the frictional power. When the friction shear stress $\mu p$ is proportional to the pressure $p$, the area of real contact is usually proportional to $p$, so the statement that the wear rate proportional to the frictional work is often equivalent to the statement that the wear rate is proportional to the area of real contact (see below).
@@ 2 General (mainly for metals)
In 1953, J.F. Archard [5] published his wear model, which stated a proportionality between the volume of wear $V$ produced during sliding and the applied load $F_{0}$, the area of real contact, and the sliding distance $L$:
@@
where $\sigma_{\rm P}$ is the yield stress or penetration hardness. This model is referred to as Archard’s wear model which can also be written as
@@
In a later paper, Archard and Hirst [6] presented the results of an extensive series of experiments on the sliding wear of metals and concluded that when steady-state (kinetic equilibrium) surface conditions are attained, the wear rates of materials are indeed proportional to the applied load (unless a change in load causes the surface conditions to change) and independent of the apparent area of contact. These observations are expected if the area of real contact is proportional to the load, as would be the case if all contact regions deform plastically. However, in many applications, surfaces make repeated contact with other surfaces, and after many contacts one expects only elastic deformations of the surface asperities. Furthermore, for some materials like rubber or gel, no plastic deformation is expected.
@@
The transfer of material from one surface to another during sliding has been the subject of numerous studies[7]. By the late 1940s, radioactive isotopes were becoming more readily available to researchers, making it possible to study the processes of material transfer in the wear of metals. These techniques were applied by Rabinowicz [8] in his studies of metal wear. The transfer of metals in sliding contacts is extremely important in the performance of switches, relays, edgeboard connectors, and other electromechanical components.
Several microscopic mechanisms for the origin of wear have been presented. Nam P. Suh developed what is known as the delamination theory of wear [9]. Published in 1973, this fatigue-based sliding wear model involved the nucleation and propagation of subsurface cracks to form platelets, which eventually detached from the wear surface as flat debris particles.
@@
Fretting wear, which typically occurs in clamped or bolted mechanical joints, stacks of objects in transport, and electrical connectors in vibrating machinery, involves small-amplitude oscillations. It differs from reciprocating sliding wear in the trapping of debris within the contact and also because the outer region of a fretting contact may sometimes be slipping while the central portion may remain stuck. Cattaneo [21] and Mindlin [22] published a famous paper providing an elastic contact analysis of the fretting problem.
@@
Fretting wear may also involve corrosion and the formation of oxidized debris, hence the term fretting corrosion. Corrosion may be facilitated by water capillary bridges, which form spontaneously between closely spaced hydrophilic surfaces in a humid atmosphere.
[…]
@@
Erosion of solids in liquid can occur when cavitation bubbles collapse [25]. Cavitation bubbles form in a liquid when the pressure drops locally below the saturated vapor pressure $p_{\rm sat}$. The bubbles grow under low ambient pressure and collapse when the surrounding liquid pressure increases again above $p_{\rm sat}$. Bubble collapse near solid walls can generate high-velocity liquid jets and shock waves that impose intense pressure loads on the wall. These pressure loads are responsible for erosive damage on solid surfaces, as observed in applications such as liquid fuel injection, hydrodynamic power generation, and marine propulsion. Conversely, the pressure loads from collapsing bubbles are exploited in applications such as shock wave lithotripsy, drug delivery, and surface cleaning.
Many forms of wear exhibit transitions, particularly during the early stages of contact (run-in). For example, in solid particle impingement and cavitation erosion, several damage stages are commonly identified: an incubation period with negligible erosion rate, an acceleration period during which the erosion rate increases to a maximum, a deceleration (attenuation) period, and a terminal period where a final steady state is reached. These stages arise from phenomena such as initial surface conditioning, cold working and hardening, and the embedment of impinging particles.
To summarize, key concepts in understanding wear include adhesion, the true area of contact, interfacial material transfer, critical angles for maximum abrasion and erosion, and the mechanics of rough surface contact.
3 Rubber wear
@@
When a needle is pressed into rubber, it creates a funnel-shaped depression. If the applied force is small, this deformation may be primarily viscoelastic and not involve bond breaking. Schallamach showed that as the needle moves, its leading edge contacts the sidewall of the funnel, causing the rubber to bulge in front of the needle [see Fig. 7(b)]. The elastic force acting on the needle has a vertical component directed upward, causing the needle to lift the rubber adhering to it. The bulge forms a lip. As the needle moves over the lip, the rubber retracts, corresponding to the tip jumping forward on the surface and producing the stick-slip motion and force fluctuations shown in Fig. 6(c) (top).
@@
Glancing incidence— Here we first present the basic concepts for the case of glancing impact, where the mechanism of material removal has been found to be very similar to that occurring during sliding abrasion by a blade.
Arnoldt and Hutchings [35] found that during erosion at glancing impact, a series of ridges running transversely to the impact direction is produced during the initial run-in stages of erosion, just as in scraping with a razor blade. This phenomenon is illustrated in Fig. 8. Impacting particles slide over the surface and deform the ridges, causing the growth of fatigue cracks from the base of each ridge, as shown in Fig. 11(a). As in the case of abrasion by a blade, it is assumed that the growth of these fatigue cracks is the rate-determining step in the erosive wear process.
Abrasion by a blade has been successfully modeled by Southern and Thomas [27], and their model was adopted by Arnoldt and Hutchings for a theory of erosive wear. To extend this model to erosive wear, it is necessary to estimate the friction force imposed by each impacting particle on the surface, and the distance over which the particle slides during contact. For blade abrasion, these factors can be controlled or measured, but in erosion, they depend on impact dynamics.
@@
To apply this equation, one needs to determine the force $F$ a particle exerts on a ridge, as well as the width $w$ and the length $L$ over which the particle slides on the ridge. They estimated $F$ and $L$ by solving Newton’s equations for an impacting particle, assuming that the normal force can be derived from a rigid flat punch model for the particle-solid interaction, and that the tangential (frictional) force is given by $F=\mu F_{0}$, where $\mu$ is the friction coefficient and $F_{0}(t)$ the normal force. For the contact width, they assumed that the normal force is high enough that $w\approx 2R$, where $R$ is the particle radius. Using this simple theory along with the measured $\Delta x(G)$ relation for several rubber types (see Fig. 12), they were able to estimate the dependence of erosion rate on angle of incidence, impact velocity, particle size, and rubber elastic modulus, all in semi-quantitative agreement with experimental data. The predicted absolute values of the erosion rate agreed with experiments within a factor of $\sim 10$.
@@
Smear film— In contrast to abrasion on sharp tracks, the rate of abrasion on blunt tracks such as concrete or well-worn grinding wheels is very sensitive to the presence of antioxidants and to the surrounding atmosphere (e.g., oxygen or nitrogen). These factors influence abrasion in a manner similar to their influence on fatigue [37].
@@
However, unlike Rabinowicz’s assumption, cracks usually propagate incrementally. The crack growth per stress cycle is $\Delta x$, which depends on the tearing energy amplitude $G$ (see Figs. 12, 16). Thus, multiple asperity contacts are generally needed to remove a particle, except when the energy release rate reaches the ultimate tear strength $G_{\rm c}$, allowing single-contact detachment (see below). This is explicitly included in the model in Ref. [12].
@@ 4 Summary and outlook
We have shown that for elastic contact without adhesion, in most cases $A/A_{0}$ is proportional to the normal force. This proportionality holds even when adhesion increases the contact area, assuming that adhesion does not manifest itself as a pull-off force, which is satisfied in most practical applications, e.g., negligible adhesion when lifting a bottle from a table [59]. However, the pressure region over which this proportionality holds is reduced by adhesion [59].

### v2

@@ 1 Introduction
Although often viewed as a technical issue, wear has posed challenges to humans for thousands of years. Even back in the Stone Age, hunters needed to consider the importance of maintaining the sharpness of their stone axes or knives. Nowadays, wear is not only a major reason for the failure of mechanical devices, but also a contributor to air pollution. In recent decades, environmental concerns related to wear have received increasing attention. Particles released from wear processes, such as those from tires or brakes, pose significant health risks. Particles with aerodynamic diameters greater than $\sim 10\ {\rm\mu m}$ are typically deposited in the nose or throat and cannot penetrate the lower tissues of the respiratory tract. However, smaller particles can enter the lungs and translocate into vital organs due to their size, causing significant human health consequences. The degree of toxicity becomes greater for smaller particles.
@@
Several comprehensive reviews on wear are available in the literature. For instance, Blau [1] focuses mainly on metals, while Muhr and Robbins [2] and Veith [3] have investigated rubber wear. A broader historical perspective on wear and tribology can be found in the beautiful book by Dowson [4].
In this article, we focus on sliding wear but also briefly discuss fretting (a form of sliding wear) and erosion, where the surface of a solid is exposed to a beam of impacting hard particles. We begin by reviewing the development of wear theories, primarily in metals. Then, we shift focus to rubber-like materials, highlighting the most influential works by pioneers such as Schallamach, Grosch, Southern and Thomas, and Arnold and Hutchings. We also consider recent studies that attempt to include the multiscale nature of roughness on real surfaces.
Figure 1: Wear in a brake system. A solid bar (pin) is pressed against a rotating cylinder (disc). If the rotating disc does not wear, then the wear of the pin is equivalent to the disc moving toward the pin. The thickness of the removed material can be obtained as the difference in circumferences between two displaced discs. The pressure is assumed to act orthogonally to the rotating disc, and the thickness of the removed material in this direction is given by $d(\theta)=d_{0}\cos\theta$. Hence, if the wear rate is proportional to the pressure, then $p\sim d(\theta)\sim\cos\theta$.
2 Foundations of Wear Theories
One of the first scientific studies of wear was the work by K.T. Reye [5]. He proposed in 1860 that the wear is proportional to the frictional work. This is in general not true since adding a monolayer of a boundary lubricant will usually change the frictional work and the wear differently [6]. In addition, the friction is usually due mainly to energy transfer into thermal movements and not due to wear. Still, the wear rate may be proportional to the (local) nominal contact pressure $p({\bf x})$, at least if frictional heating is unimportant, and since the frictional power often is proportional to the nominal contact pressure, $\mu pv$, if the friction coefficient is assumed constant then the local wear rate may be proportional to the frictional power. When the friction shear stress $\mu p$ is proportional to the pressure $p$, the area of real contact is usually proportional to $p$, so the statement that the wear rate is proportional to the frictional work is often equivalent to the statement that the wear rate is proportional to the area of real contact (see below).
@@ 2 Foundations of Wear Theories
In 1953, J.F. Archard [7] proposed a wear model establishing a proportionality between the volume of wear $V$ produced during sliding, the applied load $F_{0}$, the area of real contact, and the sliding distance $L$:
@@
where $\sigma_{\rm P}$ is the yield stress or penetration hardness. This model is now known as Archard’s wear model which can also be written as
@@
In a subsequent study, Archard and Hirst [8] reported comprehensive experimental on the sliding wear of metals. They concluded that under steady-state surface conditions (which correspond to kinetic equilibrium), the wear rates of materials are indeed proportional to the applied load (unless a change in load causes the surface conditions to change) and independent of the apparent area of contact. These observations are expected if the area of real contact is proportional to the load, as would be the case if all contact regions deform plastically. However, in many applications, surfaces make repeated contact with other surfaces, and after many contacts one expects only elastic deformations of the surface asperities. Furthermore, for some materials like rubber or gel, no plastic deformation is expected.
@@
The transfer of material between two surfaces during sliding has been intensively studied [9]. By the late 1940s, radioactive isotopes were becoming more readily available to researchers, making it possible to study the processes of material transfer in the wear of metals. These techniques were applied by Rabinowicz [10] in his studies of metal wear. Experimental studies have shown that the wear of metals often proceeds through multiple stages. Initially, metal fragments are transferred from one surface to another. These fragments are then smeared across the surface, forming a transfer layer that is typically harder than the original material. This layer subsequently undergoes oxidation, a process accelerated by the elevated temperatures generated through frictional heating. As the oxidation progresses, the bond between the transfer layer and the underlying surface becomes weakened. Eventually, the oxidized material is removed in the form of large oxide particles, which constitute the wear debris [9]. Wear particles can remain trapped between the surfaces and reattach several times before finally emerging outside of the contact. The transfer of metals in sliding contacts is extremely important in the performance of switches, relays, edgeboard connectors, and other electromechanical components.
Several microscopic mechanisms for the origin of wear have been presented. Nam P. Suh proposed the delamination theory of wear [11], a sliding wear model introduced in 1973 that is based on fatigue mechanisms. This model describes how subsurface cracks initiate and grow, leading to the formation of flat wear particles that eventually separate from the surface.
@@
Fretting wear, typically arises in situations such as clamped or bolted mechanical joints, stacks of components during transport, and electrical connectors within vibrating equipment. It is characterized by small amplitude oscillatory motion. Unlike reciprocating sliding wear, fretting wear often involves the trapping of debris within the contact region. Additionally, the outer areas of the contact may experience slipping, whereas the central region can remain adhered. Cattaneo [23] and Mindlin [24] published a famous paper providing an elastic contact analysis of the fretting problem.
@@
Fretting wear may also involve corrosion and the oxidization of debris, which is known as fretting corrosion. Corrosion may be facilitated by water capillary bridges, which form spontaneously between closely spaced hydrophilic surfaces in a humid atmosphere. During the 1960s, there was increasing interest in understanding the mechanisms of erosive wear in applications such as electrical power plants, mining machinery, and jet engines. Foundational work by Finnie in 1958 and 1960 provided key insights into the erosion behavior of ductile metals subjected to hard particle impingement [25, 26]. The main parameters governing this erosion process include particle velocity, angle of incidence, and particle flux. For ductile metals, the erosion rate was typically observed to peak at incidence angles between approximately $15^{\circ}$ and $30^{\circ}$. It was also demonstrated that the erosion rate, defined as mass loss per unit mass of impacting particles, could be modeled as a constant multiplied by the particle velocity raised to a power $n$, with $n$ ranging from about $2$ to $2.5$ for metallic materials. Observations revealed the embedment of hard eroding particles such as alumina into the surface, the development of subsurface damage gradients, and fatigue-like dislocation structures within the affected regions of the material.
@@
Erosion of solid materials in liquid environments can result from the collapse of cavitation bubbles [27]. These bubbles form when the local liquid pressure falls below the saturated vapor pressure $p_{\rm sat}$. Under low ambient pressure, the bubbles expand, and they collapse once the surrounding liquid pressure rises above $p_{\rm sat}$. When bubble collapse occurs near a solid boundary, it can produce high-velocity liquid jets and shock waves that exert extreme pressure on the surface. These transient pressure loads are responsible for causing erosive damage to solids, as seen in systems such as liquid fuel injectors, hydropower equipment, and marine propulsion units. In contrast, the same pressure effects generated by bubble collapse are intentionally utilized in applications like shock wave lithotripsy, targeted drug delivery, and precision surface cleaning.
Many sliding contacts exhibit strongly varying friction and wear rates as a function of the sliding distance, but tend to approach a steady state after sufficiently long sliding distances (see Fig. 6). This phenomenon is referred to as run-in. Achieving a steady-state after an adequate run-in period is crucial in nearly all applications involving repeated use of mechanical components, such as the piston–cylinder contact in combustion engines or tire–road interactions.
Run-in typically involves the smoothing of asperities through fracture or plastic deformation. In the case of metals, it may also include material transfer between the contacting surfaces and the formation of modified surface layers.
Figure 6: Friction coefficient as a function of the number of cycles for (metal-doped) amorphous diamond-like carbon coatings: TiC/a:C (dashed line), TiC/a:C-H (dotted line), and WC/a:C-H (solid line) sliding against steel in dry air. Adapted from [28].
In summary, fundamental concepts essential to understanding wear phenomena include adhesion, the true contact area, interfacial material transfer, critical angles associated with maximum abrasion and erosion, and the mechanical behavior of rough surface interactions.
3 Mechanisms of Rubber Wear
@@
When a needle is pressed into rubber, it creates a funnel-shaped depression. If the applied force is small, this deformation may be primarily viscoelastic and not involve bond breaking. Schallamach showed that as the needle moves, its leading edge contacts the sidewall of the funnel, causing the rubber to bulge in front of the needle [see Fig. 8(b)]. The elastic force acting on the needle has a vertical component directed upward, causing the needle to lift the rubber adhering to it. The bulge forms a lip. As the needle moves over the lip, the rubber retracts, corresponding to the tip jumping forward on the surface, producing the stick-slip motion and force fluctuations shown in Fig. 7(c) (top).
@@
Glancing incidence—Here we first present the basic concepts for the glancing impact, a case in which the mechanism of material removal is similar to that observed in sliding abrasion by a blade.
Arnoldt and Hutchings [38] observed that, during erosion under glancing impact conditions, a series of ridges oriented transverse to the impact direction is formed during the initial run-in phase. This pattern is similar to that produced by scraping with a razor blade and is illustrated in Fig. 9. As the particles impact and slide along the surface, the ridges get deformed, leading to the propagation of fatigue cracks from the base of each ridge, as shown in Fig. 12(a). Similar to blade-induced abrasion, the propagation of these fatigue cracks is considered to be the rate-determining step in the erosive wear process.
They adapted the model developed by Southern and Thomas [30] to describe blade abrasion. In order to extend this model to the case of erosion, it is necessary to estimate the frictional force exerted by each impacting particle on the surface and the distance over which the particle slides during the contact time. While these parameters can be experimentally controlled or measured in blade abrasion, they are governed by the dynamics of impact in erosive conditions.
@@
To apply this equation, one needs to determine the force $F$ a particle exerts on a ridge, as well as the width $w$ and the length $L$ over which the particle slides on the ridge. They estimated $F$ and $L$ by solving Newton’s equations for an impacting particle, assuming that the normal force can be derived from a rigid flat punch model for the particle-solid interaction and that the tangential (frictional) force is given by $F=\mu F_{0}$, where $\mu$ is the friction coefficient and $F_{0}(t)$ the normal force. For the contact width, they assumed that the normal force is high enough that $w\approx 2R$, where $R$ is the particle radius. Using this simple theory along with the measured $\Delta x(G)$ relation for several rubber types (see Fig. 13), they were able to estimate the dependence of erosion rate on the angle of incidence, impact velocity, particle size, and rubber elastic modulus. These dependencies are all in semi-quantitative agreement with experimental data. The predicted absolute values of the erosion rate agreed with experiments within a factor of $\sim 10$.
@@
Smear film—In contrast to abrasion occurring on sharp tracks, the abrasion rate on blunt tracks, such as concrete or worn grinding wheels, is highly sensitive to the presence of antioxidants and to the surrounding atmosphere (e.g., oxygen or nitrogen). These factors influence the abrasion process in a way that is comparable to their effects on fatigue behavior [40].
@@
It is well known that, under specific conditions, rubber surfaces can become tacky during abrasion experiments, tire drum testing, and occasionally during real road tire usage. One proposed explanation is that thermomechanical degradation of the polymer leads to the formation of lower molecular weight species. In this context, the degradation is believed to result from stress-augmented thermal activation bond breaking [41].
@@
However, unlike Rabinowicz’s assumption, cracks usually propagate incrementally. The crack growth per stress cycle is $\Delta x$, which depends on the tearing energy amplitude $G$ (see Figs. 13, 17). Thus, multiple asperity contacts are generally needed to remove a particle, except when the energy release rate reaches the ultimate tear strength $G_{\rm c}$, allowing single-contact detachment (see below). This is explicitly included in the model in Ref. [14].
@@ 4 Summary and Outlook
These insights underline the need for future studies that bridge the gap between continuum-scale models and the molecular-scale mechanisms underlying wear in soft materials. In particular, understanding how energy dissipation mechanisms vary with length scale and how they influence crack initiation and propagation remains a key challenge. Studies of the nucleation of cracks at surfaces and measurements of the tearing energy relation $\Delta x(G)$ for micrometer-sized cracks at surfaces (which may be possible using optical microscopy) are needed for a better understanding of rubber wear processes. Investigations into the distribution of cracks on rubber surfaces after run-in are also important and may be possible using scanning electron microscopy on bent rubber samples that open up surface cracks. Moreover, the development of multiscale simulation frameworks that integrate atomistic models (e.g., molecular dynamics) with continuum mechanics will be essential to predict wear behavior under realistic conditions. Addressing these challenges will not only deepen our understanding of rubber wear but also support the design of more durable and environmentally friendly materials.
@@
We have shown that for elastic contact without adhesion, in most cases $A/A_{0}$ is proportional to the normal force. This proportionality holds even when adhesion increases the contact area, assuming that adhesion does not manifest itself as a pull-off force, which is a condition that is satisfied in most practical applications, e.g., negligible adhesion when lifting a bottle from a table [62]. However, the pressure region over which this proportionality holds is reduced by adhesion [62].